OpenStudy (anonymous):

If each angle in a pentagon is the same measure, what is that measure? 36° 540° 180° 108°

OpenStudy (anonymous):

Formula for finding interior angle is: \[\large \color{green}{Interior \; Angle = \frac{n-2}{n} \times 180}\] Put n = 5 here and tell me what you get..

OpenStudy (anonymous):

Put n = 5 here and solve..

OpenStudy (anonymous):

108 degrees

OpenStudy (anonymous):

The sum of the measures of the interior angles of any polygon can be calculated using the formula (n-2)180, where n = the number of sides. If the pentagon is a regular pentagon, then all of the interior angles are congruent (i.e. : 108 degrees).

OpenStudy (anonymous):

Yes absolutely you are right..

OpenStudy (anonymous):

i seen that on yahoo qibtiya

OpenStudy (anonymous):

So you got your answer @beezz12 ??

OpenStudy (anonymous):

yepp